Solve systeur xc²-3 y=0 of Eg. 2 y²-3x=0
Answer #1

о 13 -3 го 92 -3=0 23 39 го 31 = 1 from equation -322 о а) - 3 352 о -3 го 9 94 - 9 + Э го 9 Ә(33 - 3) = 0 (9-3) (81 +38 +s )t! 8h - 3 уг 1 then ور اس у го 9 ет, я з 3 у - 3 - 8 = -3+ зі (2+ 5-різ? 1 3 – 9 – 18 з і не — 10 — 183 і ус 4- у е –9 – 93 22 Kas+svi i th เg- - ) S-5- - - รัฐ 10 LVL +t- LCM ฯe ? O K AC